I don't understand how to do this equation. A friend gave me the answer, but didn't explain the work.Help.
x^2-11x+28=0

Answer by smik(40)

Okay, with this equation it's all about factoring. To factor a quadratic that hasn't got a coefficient in front of x^2, we basically just ask ourselves: what two numbers added together give us (-11) and give us 28 when multiplied.

Hmm ... don't (-7) and (-4) add to (-11) and multiply to 28??

Yes, they do.

So we get the following:

(X-7)(X-4) = 0

We're looking for all instances when y = 0. Therefore, we can deduct that:

X = 7, and X = 4

As (7-7)(7-4) = 0 >> (0)(3) = 0 >> 0 = 0
And (4-7)(4-4) = 0 >> (-3)(0) = 0 >> 0 = 0